Mir Status Report 2 April 1997

Mir 23 / NASA 4 Daily Status Reports (Edited) // Period of April 2-3, 1997
Date: 2 April 1997 / Mission Day 81

Experiment Hardware:

1. EDLS - Jerry saw a green light flash inside the drive door and on panel every three seconds regardless of activity. We have not pursued this report as a high priority item. At the present time it is believed that all of the MiSDE PCMCIA disks are full. We are waiting to hear Jerry's report on the disk status before determining whether any future EDLS sessions will be scheduled.

2. Orientation - It has been reported by the crew the Orientation session ran into difficulties when Jerry was unable to activate the system. This is still being investigated by the Russian specialists; however, it appears that it is possible, repeat, possible, radiation has hit the PCMCIA cards. They are believed to be static PCMCIA cards of the type originally used by MIPS.

Video System - Video downlink of the CNN Interview was accomplished. The video quality was poor; according to the Russian specialist, the signal was distorted due to problems between Russian ground sites.

Mir Systems

1. O2 Candles - The crew continues to burn three candles per day. A crew call down on 1 April gave a status of candles remaining onboard Mir. One candle was used from a 1989 crate (crate #10594), which was opened on 1 April. This is the fifth crate used with six crates remaining. Two bags of approximately 20 candles have been moved to Priroda. Approximately 20 loose candles are behind panel 408 in Kvant. The crate which supplied the "fire causing candle" remains with less than 20 candles.

2. Condensate Recovery System - The crew will perform an IFM on the Condensate Recovery System today. A hose will be installed between the separation unit and the purification filter to bypass a clogged line. The old purification filter will remain in the system for one month following the repair of the system. The crew continues to drink CWC water.

3. Thermal Control System (VGK) - On 27 March, the crew detected ethylene glycol on the connectors and pipes of the VGK thermal loop in Kvant II. Since telemetry did not indicate a leak in the system, experts estimated the amount of fluid leak was below the threshold for detection (50-60 ml). It was recommended the crew dry the affected area and monitor for the return of the liquid. Since initial detection, the crew has provided video downlink of the leak and reported finding approximately 100 ml of coolant at the leak site. The leakage rate is estimated to be 8-9 ml/hr. With increasing levels of ethylene glycol vapor, it will become irritating to a crewmember before it becomes detrimental to the crew's health. To date, a total of four tubes of sealant have been found on board to repair this loop, which will be performed 2 April. After applying sealant, the internal loop pressure will be decreased to create a patch. Subsequently (on 4 April) the loop pressure will be increased to the nominal pressure. As a result of this loop being turned off, the crew reported it was "...getting pretty hot in Kvant II". Attitude maneuvers to cool Kvant II were performed, while taking precautions not to create thermal problems in Kvant. On 2 April, the crew discovered two areas of corrosion. These areas are clearly the point of leakage. The crew applied sealant on the areas are trying to keep the delta pressure constant. The ground informed the crew that it will probably get a little warmer on the Mir.

4. Thermal Control System (KOB-1/2) - There is also concern of leakage in coolant loop KOB-1. According to V. Blagov, the leak is in the bellows and cannot leak into the habitable cabin. The pressure in the redundant loop, KOB-2, has dropped. The crew has been instructed to deactivate this loop.

5. Core Module - The crew provided video of wall dimples in the transfer node of approximately 0.5 mm in depth. The SFD expressed no concerns since the required thickness of the wall for a delta pressure of 1 atm is about 1 mm. Specialists are considering coating the wall to prevent future oxidation of the aluminum wall.

6. The crew was told to expect higher temperatures in the Base Block for the next few days. This is due to attitude changes performed to improve the temperatures in Kvant 2. The current temperature (16:00 DMT) is +31.26 deg C.

7. Progress - Mr. Blagov said Progress is on schedule for a 6 April launch. He said that the new Altair transmitter was not ready in time for the loading of the Progress.

8. Altair - On March 27 the Altair began its 40 day journey from east to west. The satellite should complete the move by 10 May, in time to support STS-84.

Environmental Parameters (Information from 1st April)

Pressure (mmHg) Base Block 750 752
ppO2 (mmHg) 157 161
ppCO2 (mmHg) 4.4 4.7

-----------------------------------------------
Date: 3 April 1997 / Mission Day 82

Mir Station System/Subsystem Activities
Replace PTAB Unit #2 in Priroda
Clean Mir power connectors x-301 and x-302

Mir Systems

Mr. Blagov reported the Vosdukh (CO2 Removal System) was shutdown due to a lack of thermal control from the VGK system in the core module. Some time ago, the core module's VGK and KOH-1V loops (which were originally independent loops) were re-plumbed to create one loop. This was done to reduce the number of required pump panels (one pump panel can support both loops). Earlier in the week, the pressure of the KOH-1V began dropping. Today, the delta pressure at the pump panel dropped to zero, resulting in a shutdown of the loop. The plans are to separate the loops back to two independent loops and determine if another leak exists and isolate it. Mr. Blagov estimated the on-board supply of LiOH canisters would last 10 days, and stated a 15 day supply is manifested on the Progress, docking next week. Currently, the priority of the loop repairs has not been established.

Note: This VGK loop is a separate loop from the one discussed in earlier reports. At last report, the other VGK loop (in Kvant 2) was being patched with an IFM procedure the ground had sent up, and was nearly complete.

Leak checks are in progress on the Kvant 1 VGK.

Last past news: No leaks were found in the VGK. The crew was asked to reconfigure the loop back to the Core Module. The Crew agreed to do this, but requested further work on leak checking be schedule for tomorrow.

Environmental Parameters (Information from 2nd April)
Parameter Min. Max.
Pressure (mmHg) Base Block 748 751
ppO2 (mmHg) 157 159
ppCO2 (mmHg) 4.7 4.9

Planning for MIR23 crew:
MD 83: IFM of VGK Coolant Loop in Kvant 2;
Disassemble Kristall Solar Array Functional Test Setup

MD 84: Additional Time (if required) for IFM of VGK Coolant Loop in Kvant 2
Wet Cleanup; Condensate Presence Monitoring Onboard; O2 Candle Burning

  • submit to reddit


Loading



Humans to Mars Summit 2019
Military Space Situational Awareness April 1 - April 2, 2019, London, UK
MilSatCom USA conference will return to Washington from the 26th – 27th June 2019
Space Access Conference 2019
Mission To Methonē by Les Johnson - Baen Books




Join our mailing list




Monthly Archives

About this Entry

This page contains a single entry by Keith Cowing published on April 2, 1997 9:10 PM.

Mir Status Report 1 April 1997 was the previous entry in this blog.

Mir Status Report 3 April 1997 is the next entry in this blog.

Find recent content on the main index or look in the archives to find all content.